Calculus II, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 06

1. Section 14.2: Problems 41, 43, 49, 51, 57, 61, 63. Hint for problems 61, 63: Read ‘Changing to Polar Coordinates’ on page 781.

2. Section 14.3: Problems 19, 21, 53, 60, 65, 67, 69, 81, 91.